### v2

# Motivating Next-Gen Accelerators with Flexible $N:M$ Activation Sparsity via Benchmarking Lightweight Post-Training Sparsification Approaches

Abstract
[…]
Motivating Next-Gen Accelerators with Flexible $N:M$ Activation Sparsity via Benchmarking Lightweight Post-Training Sparsification Approaches
Shirin Alanova 1, Kristina Kazistova 1, Ekaterina Galaeva 1, Alina Kostromina 1, Vladimir Smirnov 1, Dmitry Redko 1, Alexey Dontsov 1, Maxim Zhelnin 1, Evgeny Burnaev 1,2, Egor Shvetsov 1 1 Skolkovo Institute of Science and Technology 2 Artificial Intelligence Research Institute 3 HSE University indicates equal contribution.
1 Introduction
Large Language Models (LLMs) have intensified demand for efficient inference. A common rule of thumb suggests serving speed for a dense $N$-parameter model scales as $\propto 1/\sqrt{N}$ (Erdil, 2025). Inference is often accelerated via quantization (Frantar et al., 2022; Shvetsov et al., 2024) or sparsification (Frantar and Alistarh, 2023; Maximov et al., 2025), with sparsity reducing both compute and memory I/O.
Weights vs. Activations. Although weight and activation sparsity yield the same theoretical FLOP count, they differ in practice. Weight sparsity enables static compression but can irreversibly degrade model quality, whereas activation sparsity is input-adaptive and better preserves model capacity though it requires an additional pruning step for each input.
Accelerating LLMs with Sparse Activations. Naturally, pruning is most effective when the values to be pruned already have low magnitudes or are zero, which is often the case for some LLMs’ intermediate representations (Liu et al., 2023; Li et al., 2022). Activation sparsity was employed to accelerate the decoding stage by enabling faster sparse vectordense matrix multiplications up to 2× speedup using specialized kernels (Song et al., 2024b, a; Liu et al., 2024; Lee et al., 2024). These gains are most pronounced for batch size 1 and diminish as batch size increases (Shrestha et al., 2025). The speedup primarily arises because rows of the dense matrix corresponding to zero elements in the sparse vector can be skipped during computation. Moreover, these methods often require some predictive mechanism to upload required weight indices into memory ahead of time Liu et al. (2023).
Semi-structured N:M activation sparsity, keeping $N$ non-zeros per block of size $M$, can extend benefits beyond single vector–matrix products by employing hardware support. Although 2:4 activation sparsity has been explored for training (Haziza et al., 2025; Wang et al., 2024), and post-training weight sparsification Maximov et al. (2025); Han et al. (2015); Frantar and Alistarh (2023); Kurtić et al. (2023) post-training N:M activation pruning is largely unexplored because it is not supported on commercial hardware.
[…]
Focus and Motivation. The primary goal of this work is twofold: to motivate the development of hardware that natively supports semi-structured activation sparsity, and to benchmark existing approaches for inducing such sparsity in activations. Generally, these approaches address two key challenges:
(P1) Selection Strategy. As with weight sparsification, the choice of which activations to retain critically affects model accuracy (Zhelnin et al., 2025).
(P2) Error Mitigation. While post-training fine-tuning can recover lost performance, it is often impractical due to computational cost or risks to safety alignment (Kharinaev et al., 2025). We therefore mainly focus on lightweight, plug-and-play methods that require minimal (e.g., WikiText) or zero calibration data.
Our Contributions:
  • First, we demonstrate that activation sparsity outperforms weight sparsity. At matched sparsity levels, activation pruning consistently yields higher accuracy than weight pruning across four diverse large language models Llama2-7B-chat, Llama3.1-8B-Instruct, Qwen2.5-7B-Instruct and Gemma3-4B-Instruct highlighting activations as a more promising target for future sparse accelerators.
  • Second, we benchmark four plug-and-play error mitigation techniques, three of which are applied to semi-structured activation sparsity for the first time, including statistical approaches such as median shift and variance correction. We also evaluate three selection criteria. These methods establish strong, retraining-free baselines that require minimal metadata, aligning well with hardware constraints.
  • Finally, we analyze a wide range of structured sparsity patterns and show that larger patterns dramatically improve model fidelity: the 16:32 pattern achieves performance close to unstructured 50% sparsity and retains over three times more accuracy than the conventional 2:4 pattern. Nevertheless, considering implementation trade-offs, we advocate for 8:16 as the optimal balance it offers twice the accuracy retention of 2:4 while remaining highly practical for near-term hardware adoption.
Together, these results provide concrete evidence that expanding hardware support beyond static 2:4 weight sparsity can unlock significant efficiency gains without compromising model quality, thereby motivating the next generation of sparsity-aware architectures.
2 Methods
In the sections below, we formally define the pruning criteria and error mitigation strategies evaluated in this work. A brief summary of these methods is provided in Table 1 and more detailed Table is available in Appendix 6.
Table 1: Summary of evaluated pruning metrics and transformations. Full descriptions are provided in Appendix 6.
Short name Description
Pruning metrics
ACT Activation magnitude
WT Weight magnitude
CLACT* Context-aware cosine-loss-inspired
Amber-Pruner Weight-informed importance
Transformations
D-PTS* Dynamic per-token shift
S-PTS Static per-token shift
L-PTS* Learnable per-token shift
VAR* Variance correction
VAR+L-PTS* VAR + learnable shift
R-Sparse Sparse activations + low-rank weights
2.1 Preliminaries
For a linear layer with weights $\mathbf{W}$, input activations $\mathbf{X}$, and output $\mathbf{Y}$:
$\mathbf{Y}=\mathbf{X}\mathbf{W}^{\top}.$ (1)
We construct a sparsity mask $\mathbf{M}$ using a metric $S$ and threshold $t$:
$\mathbf{M}_{ij}=\begin{cases}1,&S(\mathbf{X}_{ij})\geq t,\\ 0,&S(\mathbf{X}_{ij})<t,\end{cases}$ (2)
yielding
$\mathbf{Y}_{p}=(\mathbf{X}\odot\mathbf{M})\mathbf{W}^{\top}.$ (3)
Unstructured sparsity applies a global threshold over all elements. Semi-structured N:MN{:}M sparsity partitions rows (or columns) into non-overlapping blocks of size $M$ and keeps the top-$N$ elements per block by $S$ (e.g., 2:4 removes 50%) (Hu et al., 2024). We also study 8:16, which retains the same 50% density with higher flexibility and modest metadata overhead (Maximov et al., 2025).
2.2 Pruning Criterion
ACT: This is the magnitude activation pruning metric (ACT), defined as the absolute value of the element $\mathbf{X}_{ij}:S_{\textit{ACT}}(\mathbf{X}_{ij})=|\mathbf{X}_{ij}|$.
WT: This is the weight-based pruning metric (WT), defined as the absolute value of the corresponding weight $\mathbf{W}_{ij}$: $S_{\textit{WT}}(\mathbf{W}_{ij})=|\mathbf{W}_{ij}|$.
CLACT: Inspired by output cosine similarity in (Mi et al., 2025), we propose Cosine Loss ACTivation (CLACT), a context-aware score that emphasizes activations aligned with their row/column energy:
$S_{\textit{CLACT}}(\mathbf{X}_{ij})=\dfrac{|\mathbf{X}_{ij}|}{\sqrt{\sum_{k=1}^{h}\mathbf{X}^{2}_{ik}}}\,\sqrt{\sum_{p=1}^{l}\mathbf{X}^{2}_{pj}},$ (4)
where $h$ is the hidden dimension and $l$ is the sequence length.
Amber-Pruner: Following (An et al., 2025), we (i) remove weight outliers outside the 0.5–99.5 percentiles, (ii) standardize the remaining weights, and (iii) score each activation as $S_{\textit{Amb.-Pr.}}(\mathbf{X}_{ij})=|\mathbf{X}_{ij}|\cdot\mathcal{L}(\hat{\mathbf{W}}_{:,j})$, where $\mathcal{L}(\cdot)$ is the channel-wise $\ell_{2}$ norm.
CLACT is context-aware (and for l=1l{=}1 reduces to an $\ell_{1}$-type criterion), whereas Amber-Pruner leverages weight magnitudes but is not context-aware.
2.3 Transformations for Error Mitigation
D-/S-/L-PTS (dynamic/ static/ learnable per-token shift) centers activations near zero: $\hat{\mathbf{X}}=\mathbf{X}-\bm{\eta}$, and uses the compensated form $\mathbf{Y}_{p}=((\hat{\mathbf{X}}\odot\mathbf{M})+\bm{\eta})\mathbf{W}^{\top}$ (Chua et al., 2024), where $\bm{\eta}=\overline{\mathbf{X}}$ for D-PTS; S-PTS uses a fixed $\bm{\eta}$ collected in a short warm-up; L-PTS learns $\bm{\eta}$.
VAR applies per-token variance correction after pruning: $\mathbf{Y}_{p}=\bm{\nu}(\mathbf{X}\odot\mathbf{M})\mathbf{W}^{\top}$, with
$\nu=\sqrt{\dfrac{\operatorname{Var}[\mathbf{X}]}{\operatorname{Var}[\mathbf{X}\odot\mathbf{M}]}}.$ (5)
VAR+L-PTS combines VAR scaling with the learnable shift. R-Sparse combines activation sparsity with a low-rank weight approximation (Kamirul et al., 2025) (details in Appendix B).
For methods requiring calibration/learning, we use WikiText-2: S-PTS stores a fixed bias vector $\bm{\eta}$, L-PTS (and VAR+L-PTS) learns $\bm{\eta}$, and R-Sparse learns low-rank factors; all other methods require no learned parameters.
2.4 Evaluation & Models
We evaluate the methods in Table 1 in two stages. First, we use Core Datasets (BoolQ, WinoGrande, PIQA, ARC-Easy) to screen all methods. We then focus on the most promising approaches on Qwen2.5-7B-Instruct and Llama3.1-8B-Instruct, and extend evaluation to Extended Datasets (HellaSwag, OpenBookQA, RTE, MMLU, Lambada_standard, Lambada_openai, IFEval). When calibration is required, we use WikiText-2, following common compression practice Egiazarian et al. (2024); Frantar et al. (2022); Van Baalen et al. (2024). All results are obtained with LM Eval Harness Gao et al. (2023); dataset details are in Table 9. We report results for Llama2-7B-chat, Llama3.1-8B-Instruct, Qwen2.5-7B-Instruct and Gemma3-4B-Instruct. For Qwen2.5-7B-Instruct, we do not sparsify key/query/value activations due to severe degradation observed in preliminary experiments.
Figure 1: Comparison of unstructured sparsity in activations (ACT) and weights (WT) averaged across four datasets at varying sparsity ratios. Higher is Better. More detailed results are presented in Appendix Table 10.
Figure 2: Comparison of sparsity patterns with unstructured sparsity. 50% and 70% correspond to unstructured sparsity. Lower is Better. More detailed results are presented in Appendix Table 7.
3 Results
3.1 Sparse Weights vs. Activations
In Figure 1 and Table 10, we demonstrate that unstructured weight sparsification causes greater model degradation than unstructured activation sparsification at the same sparsity levels: {20%, 50%, 70%, 90%}. For this evaluation, we specifically use unstructured magnitude-based sparsification, as it is less damaging than semi-structured sparsification and thus provides a lower bound on performance degradation.
3.2 Optimal semi-structured sparsity patterns
Our preliminary investigation demonstrates that while the 16:32 pattern achieves performance close to unstructured sparsity (a 5.4% drop versus 4.5% for 50% unstructured), it requires more metadata and greater resources for gather operations, as discussed in Section 1. Therefore, we focus on the 8:16 pattern, despite its higher performance drop of 7.38%. For comparison, the 2:4 pattern results in a 14.35% drop. These results are shown in Figure 2 and Table 7 in Appendix, we used only magnitude pruning to obtain these results. By demonstrating the superior model quality of 8:16 sparsity, our work incentivizes hardware designers to invest in or at least consider supporting the 8:16 pattern.
Table 2: Average relative performance drop (%, lower is better; negative = improvement) across four datasets, averaged over Llama2-7B-chat, Llama3.1-8B-Instruct, Qwen2.5-7B-Instruct and Gemma3-4B-Instruct. Act/Wt: activation/weight sparsity. Selection: ACT/CLACT/Amber-Pruner; transformations: VAR, D-PTS, S-PTS, L-PTS, R-SPARSE. Methods marked with an asterisk (*) are proposed in this paper. Full results: Appendix 11 and  12.
Target Pattern Method Avg drop ($\downarrow$)
Act u50 ACT $3.82$
Wt 2:4 WT $24.49$
Act 2:4 ACT $9.67$
Act 2:4 CLACT* $7.79$
Act 2:4 Amber-Pruner $7.85$
Act 2:4 VAR* $6.09$
Act 2:4 D-PTS* $5.84$
Act 2:4 S-PTS $4.29$
Act 2:4 L-PTS* $8.79$
Act 2:4 R-SPARSE (64) $7.70$
Act 2:4 R-SPARSE (128) $8.05$
Wt 8:16 WT $17.68$
Act 8:16 ACT $5.47$
Act 8:16 CLACT* $2.29$
Act 8:16 Amber-Pruner $1.56$
Act 8:16 VAR* $3.30$
Act 8:16 D-PTS* $2.07$
Act 8:16 S-PTS $0.61$
Act 8:16 L-PTS* $5.32$
Act 8:16 R-SPARSE (64) $1.52$
Act 8:16 R-SPARSE (128) $2.63$
3.3 Results on Single/Multi-choice Datasets
3.3.1 Evaluation of Pruning Criteria
We evaluate CLACT, Amber-Pruner, and magnitude pruning as a baseline. The main results for the 2:4 and 8:16 sparsity patterns are presented in Table 2. On average, both CLACT and Amber-Pruner outperform magnitude pruning by at least 2%, however, we observe no clear winner between them. As noted in Section 2.2, these criteria are designed for different purposes: CLACT adjusts based on context, while Amber-Pruner adjusts based on weight magnitudes. Notably, for Llama3.1-8B-Instruct under the 2:4 sparsity pattern, simple magnitude pruning outperforms both advanced criteria, underscoring model and architecture-specific sensitivities to pruning strategies.
3.3.2 Evaluation of Transformations
Our main results are presented in Table 2. Surprisingly, we find that simple methods such as dynamic and static per-token shifts (D-PTS, S-PTS) outperform most other approaches. The second most effective methods are VAR and R-SPARSE. We also observe that increasing the number of dimensions in R-SPARSE (from 64 to 128) leads to worse performance, which may indicate overfitting on the calibration data. Finally, we note that L-PTS, the approach with learnable per-token shifts, significantly underperforms compared to its static counterpart, S-PTS.
3.4 Instruction-Following Tasks
Table 3 presents instruction-following performance on the IFEval benchmark for Llama3.1-8B-Instruct and Qwen2.5-7B-Instruct, evaluated under two semi-structured sparsity patterns (2:4 and 8:16) and four activation transformation methods: S-PTS, D-PTS, R-Sparse, and VAR. First of all, we observe a strong model degradation on generative tasks. Second of all, we see that VAR is the strongest performer overall, especially for Llama3.1-8B-Instruct. S-PTS/D-PTS are competitive and lightweight, and R-Sparse lags significantly, particularly at 2:4. We speculate that while semi-structured patterns are good for prefill stage in LLMs they significantly degrade performance during decode stage. However, as discussed in Section 1 decode stage for single vector can be accelerated with more flexible approaches.
Table 3: Instruction-following (IFEval) prompt-level accuracy reported as PS/PL (prompt-level strict/loose accuracy).
Model Method 2:4 8:16
Llama3.1-8B ORIG 0.4455/0.4861 0.4455/0.4861
S-PTS 0.1682/0.1904 0.2995/0.3327
D-PTS 0.1941/0.2015 0.2828/0.3198
R-Sparse 0.0869/0.0979 0.2089/0.2311
VAR 0.2237/0.2458 0.3161/0.3586
Qwen2.5-7B ORIG 0.7135/0.7394 0.7135/0.7394
S-PTS 0.4325/0.5176 0.5194/0.5804
D-PTS 0.4399/0.5120 0.5434/0.5989
R-Sparse 0.2736/0.3457 0.3697/0.4196
VAR 0.4565/0.5342 0.5249/0.5896
3.5 Results for Unstructured Pruning
We evaluate D-PTS, VAR, and two selection criteria in this experiment: CLACT and Amber-Pruner using the Llama3.1-8B-Instruct model. Results in Table 4 indicate that VAR is the most effective transformation under unstructured sparsity. Moreover, CLACT outperforms Amber-Pruner by a wider margin here than in our semi-structured pruning experiments. These findings suggest two key insights: (1) No single method emerges as optimal for both unstructured and semi-structured sparsity. (2) The methods proposed in this work, VAR and CLACT, demonstrate strong generalization and are well-suited for both semi-structured and unstructured activation pruning.
Table 4: Performance comparison of pruning methods under 50% and 70% unstructured sparsity for Llama3.1-8B-Instruct model.
Method ArcE BoolQ PIQA Wino Grande (%) Avg. Drop
Original $0.821$ $0.839$ $0.800$ $0.734$
Unstructured 50%
ACT $0.777$ $0.820$ $0.771$ $0.686$ $4.450$
D-PTS $0.786$ $0.825$ $0.781$ $0.690$ $3.600$
VAR $0.784$ $0.819$ $0.776$ $0.705$ 3.470
CLACT $0.780$ $0.825$ $0.766$ $0.700$ $3.890$
Amber $0.768$ $0.820$ $0.763$ $0.702$ $4.450$
Unstructured 70%
ACT $0.558$ $0.631$ $0.647$ $0.548$ $25.320$
D-PTS $0.565$ $0.624$ $0.651$ $0.534$ $25.680$
VAR $0.614$ $0.651$ $0.676$ $0.532$ 22.660
CLACT $0.555$ $0.604$ $0.627$ $0.524$ $27.670$
Amber $0.487$ $0.594$ $0.590$ $0.539$ $30.680$
3.6 Combination of Methods
Next, we evaluated combinations of multiple approaches to explore potential performance gains. These combinations and their results are presented in Table 8. As shown, none of the evaluated combinations outperforms any single method, highlighting the challenges of naively combining them.
Table 5: Llama3.1-8B-Instruct with 8:16 activation sparsity: aggregated results across layer subsets. ORIG. AVG. = 0.6726. Drop is computed w/o perplexity. Full results: Appendix 13.
Method Layers PPL AVG. Drop $\downarrow$
LS+L-PTS all 9.6036 0.6047 10.90%
LS+L-PTS key,out,gate,down 8.3483 0.6385 5.43%
LS+L-PTS key,value,gate,down 8.0821 0.6503 3.56%
LS+L-PTS+VAR all 9.4983 0.6056 10.60%
LS+L-PTS+VAR key,out,gate,down 8.2930 0.6422 4.64%
LS+L-PTS+VAR key,value,gate,down 8.0259 0.6516 3.36%
3.7 Layer Sensitivity
We study layer sensitivity to 8:16 activation sparsity on Llama3.1-8B-Instruct using the Extended Datasets (Table 5). This analysis focuses on learnable methods, mainly L-PTS and LS (learnable diagonal scaling). Although learnable approaches underperform on average in our broader experiments, the best 8:16 configurations for Llama3.1-8B-Instruct rely on learnable parameters (Table 5). We observe that the FFN up projection and the attention out projection [^1] The output projection of the attention mechanism. It combines outputs from all attention heads and projects them back to the model’s hidden dimension. are the most sensitive: pruning them causes the largest drops, suggesting they should be preserved or handled with extra care. While this may not generalize to all layers, it indicates that layer importance under activation sparsity is highly non-uniform.
4 Discussion
The performance gap between multiple/single-choice benchmarks (e.g., BoolQ, PIQA) and IFEval likely stems from differences in the inference stages they emphasize. Core QA benchmarks primarily stress the prefill phase, whereas IFEval evaluates both prefill and autoregressive generation. Our evaluation remains valid: semi-structured patterns like 2:4 and 8:16 are especially effective at accelerating the prefill stage, which often dominates inference latency.
5 Conclusion
This work establishes that post-training activation pruning is significantly more accuracy-preserving than weight pruning in large language models. Across four diverse architectures (Llama2-7B, Llama3.1-8B, Qwen2.5-7B, and Gemma3-4B), we demonstrate that activation sparsity consistently retains model capabilities better than weight sparsity at matched sparsity levels.
Our evaluation reveals that lightweight error mitigation techniques particularly CLACT, D-PTS, and VAR establish strong, hardware friendly baselines requiring minimal calibration data. Through systematic analysis of semi-structured patterns, we find that 16:32 approaches unstructured 50% sparsity in fidelity, while 8:16 emerges as the optimal near term target.
6 Limitations
Key limitations: First, all evaluations use software emulation without hardware measurements of speedup or energy efficiency. Second, layer sensitivity analysis remains preliminary; while FFN up-projections and attention out-projections appear most vulnerable, broader architectural studies are needed. Third, generative performance (IFEval) degrades significantly more than multiple-choice QA under sparsity, revealing an evaluation bias toward prefill-dominated workloads. The anomalous improvements on Qwen2.5-7B-Instruct for some benchmarks further highlight dataset-specific artifacts rather than genuine capability preservation. Our hardware overhead analysis in Appendix A is a rough estimate and more precise analysis is required.
Appendix
Appendix A Hardware Implications and Computational Overhead Analysis
While our empirical results demonstrate significant accuracy benefits from flexible N:M activation sparsity, the practical value of these techniques depends critically on whether hardware implementations can overcome the computational overhead of dynamic sparsification. This section provides a comprehensive break-even analysis to determine the hardware conditions required for activation sparsity to deliver net performance and efficiency gains.
The fundamental challenge for activation sparsity is that the theoretical memory bandwidth reduction must overcome the overhead introduced by dynamic sparsification operations. Based on literature and performance models, we estimate break-even thresholds where benefits begin to outweigh costs.
A.1 Energy-Delay Product Analysis
The Energy-Delay Product (EDP) provides a comprehensive metric for evaluating whether activation sparsity delivers net efficiency benefits. For 8:16 sparsity to be worthwhile from an EDP perspective, it must overcome both computational overhead and energy costs of the sparsification process itself.
We model EDP improvement for semi-structured sparsity patterns as:
$\text{EDP}_{\text{improvement}}=\frac{\text{EDP}_{\text{dense}}}{\text{EDP}_{\text{sparse}}}\approx\frac{r\cdot\eta}{1+\alpha}$
Where:
  • $r=2.0$ is the theoretical bandwidth reduction ratio for 8:16 sparsity
  • $\eta=0.85$ is the hardware utilization efficiency (representative of practical implementations)
  • $\alpha=0.3$ is the overhead factor from sparsification operations
This overhead factor $\alpha$ is calibrated from real measurements Fang et al. (2024) demonstrated that dynamic activation sparsification incurs 30–35% additional latency on current hardware without native support, broken down as:
  • Activation magnitude computation and block-wise selection
  • Mask application and metadata handling
  • Error mitigation techniques (D-PTS, VAR), which is not included in Fang et al. (2024).
Solving for the minimum hardware acceleration factor $k$ required for net EDP benefits:
$r\cdot\eta>k\cdot(1+\alpha)$
$2.0\cdot 0.85>k\cdot(1+0.3)$
$k>\frac{1.7}{1.3}\approx 1.31$
However, due to our imprecise estimations we will consider a higher amortized $k>1.6\times$ required for speedup.
A.2 Hardware Implementation Requirements
[…]
  • Dedicated sparsity controllers: On-chip circuitry that can generate sparsity masks with minimal latency, reducing the 15–20% selection overhead
  • Hardware-supported statistical units: Specialized units for variance/mean calculation required by error mitigation techniques, eliminating their computational overhead
  • Hierarchical sparsity support: Different patterns for different layer types based on sensitivity analysis (Section 3.7)
  • Bandwidth-optimized gather operations: Specialized memory controllers that maintain high efficiency despite irregular access patterns
Recent hardware developments show promising directions: NVIDIA’s Blackwell architecture includes a dedicated hardware decompression engine Jarmusch and Chandrasekaran (2025), while the LazyGPU microarchitecture enables lazy memory request issuing Liu et al. (2025).
[…]
Table 6: Brief description of evaluated activation pruning metrics (top) and transformations (bottom). Abbreviations in bold with an asterisk (*) denote methods proposed here or first evaluated with sparse activations.
Short Name Method Key Mechanism
Pruning metrics
ACT Magnitude Pruning Selects based on activation magnitude
WT Weight-based Pruning Selects by corresponding weight magnitude
CLACT* Cosine Loss Activation A metric inspired by cosine similarity from Mi et al. (2025)
Amber-Pruner (2025) Weights Importance A metric which accounts for important weights after outlier removal and normalization
Transformations
D-PTS* Dynamic Per-Token Shift Batch-wise dynamic centering of activations before sparsification
S-PTS (2024) Static Per-Token Shift Fixed centering of activations before sparsification using a per-token bias value pre-collected on WikiText-2
L-PTS* Learnable Per-Token Shift Fixed centering of activations before sparsification using per-token bias value learned on WikiText-2
VAR* Variance Correction Token-wise variance normalization after sparsification
VAR+L-PTS* Scaling + Learnable Shift Apply VAR scaling, then add per-token bias value learned on WikiText-2
R-Sparse (2025) Rank-Aware Sparsity Combines sparse activations with weight low-rank SVD factors learned on WikiText-2
Appendix B R-Sparse Details
Finally, we include R-Sparse (Kamirul et al., 2025), which combines activation sparsity with a low-rank approximation of the weight matrix. Instead of pruning solely by magnitude, R-Sparse decomposes the computation into two parts: (i) sparse channels with high-magnitude activations, and (ii) a low-rank component obtained via SVD of $\mathbf{W}$ that approximates the contribution of pruned activations.
Formally, the linear layer
$\mathbf{Y}=\mathbf{X}\mathbf{W}^{\top}$ (6)
is approximated as
$\mathbf{Y}\approx\mathbf{Y}_{s}+\mathbf{Y}_{r},$ (7)
where
$\mathbf{Y}_{s}=\sigma_{t(s)}(\mathbf{X})\mathbf{W}^{\top},$
$\mathbf{Y}_{r}=(\mathbf{X}-\sigma_{t(s)}(\mathbf{X}))(A_{r}B_{r})^{\top}.$
Here $\sigma_{t(s)}(\cdot)$ denotes sparsification of activations with threshold $t(s)$, and $A_{r}B_{r}^{\top}$ is the rank-$r$ approximation of $\mathbf{W}$ obtained from its truncated SVD. The trade-off between $\mathbf{Y}_{s}$ and $\mathbf{Y}_{r}$ is determined by a sparsity budget $s$ and rank $r$, which can be optimized via evolutionary search.
Appendix C Main extended results
Here, we present un-aggregated results. Comparisons between different sparsity patterns is presented in Table 7. In Table 11 and Table 12 we present the results of the error mitigation strategies and the selection criteria. Table 11 reports the results of the combined methods. Table 10 compares unstructured and semi-structured sparsity. Finally, in Table 13 we demonstrate results when some of the layers are excluded to evaluate layer sensitivity.
Table 7: Performance comparison of different sparsity patterns on Llama3.1-8B-Intsruct across various benchmarks. Values represent accuracy scores, with the last column showing the average performance drop relative to the original model.
ARC Easy BoolQ PIQA WinoGrande Avg Drop ($\downarrow$)
Original $0.8207$ $0.8391$ $0.8003$ $0.7340$
2:4 $0.6837$ $0.7261$ $0.7163$ $0.6110$ 14.3514.35%
4:8 $0.7272$ $0.7810$ $0.7529$ $0.6393$ 9.299.29%
8:16 $0.7525$ $0.7969$ $0.7568$ $0.6551$ 7.387.38%
16:32 $0.7698$ $0.8082$ $0.7688$ $0.6771$ 5.405.40%
50% unstructured $0.7820$ $0.8198$ $0.7714$ $0.6858$ 4.304.30%
70% unstructured $0.5580$ $0.6311$ $0.6474$ $0.5477$ 25.3225.32%
Table 8: A comparison of combined approaches with 8:16 semi-structured sparsity. Average relative performance (%) across four datasets. Values indicate performance drops (lower is better), negative values signify performance improvement. Full, non-aggregated results are available in Appendix 12.
Models
Method Llama2-7B-chat Qwen2.5-7B-Instruct Gemma3-4B-Instruct Llama3.1-8B-Instruct Average Drop ($\downarrow)$
CLACT + PTS 5.635.63% 5.06-5.06% 0.500.50% 8.558.55% 2.40%
CLACT + VAR 5.075.07% 2.90-2.90% 0.540.54% 8.598.59% 2.822.82%
Amber-Pruner + PTS 6.166.16% 3.47-3.47% 0.170.17% 7.427.42% 2.572.57%
Amber-Pruner + VAR 4.744.74% 3.63-3.63% 0.16-0.16% 8.398.39% 2.34%
L-PTS + VAR 6.876.87% 2.862.86% 3.413.41% 7.157.15% 5.075.07%
Appendix D Datasets
Detailed description of the datasets is given in Table 9.
Table 9: Datasets used to evaluate hypotheses. Prompt-level strict accuracy is the fraction of prompts for which all verifiable instructions in the prompt are followed exactly as stated. Instruction-level strict accuracy is the fraction of individual instructions that are followed exactly as stated, averaged across all instructions.
Dataset Description Metric
WikiText-2 (Merity et al., 2016) A collection of over 100 million tokens extracted from the set of verified Good and Featured articles on Wikipedia. Perplexity
ARC-Easy (Clark et al., 2018) QA benchmark for genuine grade-school level, multiple-choice science questions. The dataset contains 2251 examples for training, 570 for development and 2376 for testing. Accuracy
ARC_Challenge (Clark et al., 2018) QA benchmark for more difficult grade-school level science questions, part of the AI2 Reasoning Challenge. Designed to require deeper reasoning than ARC-Easy. Accuracy
BoolQ (Clark et al., 2019) QA benchmark for yes/no questions. The dataset contains 9427 examples for training and 3270 for testing. Accuracy
PIQA (Bisk et al., 2020) Physical commonsense QA benchmark for choosing the right answer between two options. Contains 16K train, 2K dev, and 3K test examples. Accuracy
WinoGrande (Sakaguchi et al., 2021) QA benchmark for pronoun resolution with adversarial filtering. Contains 40K train, 1267 dev, and 1767 test examples. Accuracy
HellaSwag (Zellers et al., 2019) Commonsense reasoning benchmark for sentence completion, designed to be easy for humans but hard for models. Contains 70K train and 10K validation examples. Accuracy
OpenBookQA (Mihaylov et al., 2018) Open-book question answering dataset requiring retrieval of elementary science facts. Contains 5957 4-way multiple-choice questions. Accuracy
RTE (Dagan et al., 2005; Haim et al., 2006) Recognizing Textual Entailment datasets from PASCAL challenges. Task is to classify if a hypothesis is entailed by a premise. Accuracy
MMLU (Hendrycks et al., 2020) Massive Multitask Language Understanding benchmark covering 57 subjects across STEM, humanities, and social sciences. Measures multitask accuracy. Accuracy
Lambada_Standard (Paperno et al., 2016) Word prediction task requiring broad discourse context. Target word is unpredictable from local context alone. Accuracy
Lambada_OpenAI (Paperno et al., 2016) LAMBADA test set preprocessed by OpenAI for standardized evaluation. Task remains final word prediction with long-range dependencies. Accuracy
GSM8K (Cobbe et al., 2021) Grade school math word problems requiring multi-step reasoning. Contains 7.5K train and 1.3K test examples. Accuracy (Strict) Accuracy (Flexible)
IFEval (Zhou et al., 2023) Benchmark with 541 prompts containing verifiable instructions to measure instruction-following fidelity. Accuracy (Prompt-level) Accuracy (Instruct-level)
Table 10: The performance of models with applied unstructured activation pruning. We show that even with severe sparsity (70-90%) models were able to perform decently on our benchmarks. ACT stands for activations pruning, WT — for weight pruning. OUT denotes values more than $10^{3}$, according accuracy scores most likely correspond to random.
Pruning WikiText2 $\downarrow$ ARC Easy BoolQ PIQA WinoGrande Drop $(\downarrow)\%$
Llama2-7B-chat
   Base 6.94 0.74 0.80 0.76 0.66 -
   0.2 ACT 6.96 0.74 0.80 0.77 0.66 -$\textbf{-0.33}\%$
   0.2 WT 7.49 0.72 0.80 0.76 0.66 $0.68\%$
   0.5 ACT 7.53 0.70 0.78 0.75 0.66 $\textbf{2.32}\%$
   0.5 WT 18.72 0.60 0.72 0.70 0.61 $11.10\%$
   0.7 ACT 20.11 0.56 0.64 0.65 0.53 $\textbf{19.62}\%$
   0.7 WT OUT 0.27 0.38 0.54 0.47 $43.44\%$
   0.9 ACT OUT 0.26 0.38 0.52 0.49 $43.39\%$
   0.9 WT OUT 0.27 0.38 0.53 0.48 $43.39\%$
Qwen2.5-7B-Instruct
   Base 7.46 0.69 0.86 0.75 0.60 -
   0.2 ACT 7.48 0.69 0.86 0.74 0.61 $\textbf{2.37}\%$
   0.2 WT 8.03 0.67 0.86 0.74 0.60 $3.42\%$
   0.5 ACT 8.3 0.67 0.87 0.74 0.58 $3.87\%$
   0.5 WT 43.6 0.56 0.80 0.68 0.57 $\textbf{3.42}\%$
   0.7 ACT 18.7 0.6 0.81 0.70 0.58 $\textbf{3.87}\%$
   0.7 WT OUT 0.28 0.38 0.54 0.48 $12.12\%$
   0.9 ACT OUT 0.25 0.38 0.54 0.52 $44.22\%$
   0.9 WT OUT 0.25 0.58 0.54 0.51 $\textbf{36.35}\%$
Gemma3-4B-Instruct
   Base 17.29 0.72 0.84 0.72 0.62 -
   0.2 ACT 17.60 0.71 0.84 0.72 0.60 $\textbf{3.35}\%$
   0.2 WT 18.93 0.68 0.84 0.72 0.59 $4.74\%$
   0.5 ACT 22.39 0.71 0.83 0.72 0.57 $\textbf{4.80}\%$
   0.5 WT 273 0.36 0.55 0.61 0.52 $30.89\%$
   0.7 ACT 88 0.55 0.63 0.66 0.54 $\textbf{19.57}\%$
   0.7 WT OUT 0.27 0.49 0.53 0.51 $38.81\%$
   0.9 ACT OUT 0.26 0.38 0.54 0.50 $42.64\%$
   0.9 WT OUT 0.25 0.45 0.52 0.52 $\textbf{40.60}\%$
Table 11: Semi-Structured 2:4 Sparsification - performance Metrics, for calibration, when it is required, and perplexity we use WikiText2. Average Drop is computed without accounting for perplexity.
Pruning WikiText2 $\downarrow$ ARC Easy BoolQ PIQA WinoGrande Average Drop %
   Llama2-7B-chat 6.94 0.74 0.80 0.76 0.66 -
ACT 10.23 0.66 0.71 0.71 0.60 9.43%
WT 42.40 0.57 0.65 0.69 0.56 16.52%
D-PTS 9.38 0.64 0.68 0.71 0.61 10.67%
S-PTS 9.36 0.66 0.68 0.71 0.60 10.37%
VAR 8.31 0.67 0.69 0.72 0.59 9.76%
CLACT 8.23 0.65 0.72 0.71 0.63 8.32%
Amber-Pruner 9.24 0.64 0.68 0.69 0.60 11.70%
LPTS 8.89 0.65 0.60 0.72 0.59 13.13%
LPTS + VAR 8.39 0.67 0.63 0.72 0.60 11.47%
R-SPARSE (64) 9.19 0.66 0.63 0.69 0.59 12.90%
R-SPARSE (128) 9.29 0.65 0.65 0.70 0.59 12.23%
   Llama3.1-8B-Instruct 7.21 0.82 0.84 0.80 0.73 -
ACT 16.61 0.68 0.73 0.72 0.61 14.35%
WT 20.14 0.41 0.57 0.60 0.54 33.63%
PTS 16.4 0.69 0.73 0.72 0.60 14.59%
S-PTS (N-100) 16.5 0.67 0.74 0.72 0.60 14.61%
S-PTS (N-200) 16.5 0.68 0.73 0.72 0.61 14.31%
VAR 14.17 0.70 0.73 0.73 0.62 13.11%
CLACT 19.49 0.65 0.71 0.69 0.59 17.27%
WANDA 15.86 0.66 0.74 0.69 0.61 15.01%
L-PTS 12.77 0.71 0.71 0.73 0.59 14.13%
L-PTS + VAR 12.40 0.73 0.71 0.73 0.60 13.49%
R-SPARSE (64) 15.07 0.69 0.72 0.71 0.61 15.28%
R-SPARSE (128) 16.09 0.67 0.71 0.70 0.61 16.34%
   Qwen2.5-7B-Instruct 7.46 0.69 0.86 0.75 0.60 -
ACT 10.06 0.65 0.86 0.72 0.54 4.95%
WT 35.37 0.53 0.78 0.68 0.54 12.96%
D-PTS 10.07 0.79 0.86 0.76 0.66 -6.46%
S-PTS 10.74 0.78 0.84 0.74 0.65 -4.43%
VAR 13.95 0.74 0.83 0.74 0.61 -1.48%
CLACT 11.16 0.73 0.84 0.71 0.67 -2.45%
Amber-Pruner 10.64 0.74 0.84 0.70 0.64 -1.23%
LPTS 9.13 0.67 0.81 0.72 0.58 3.66%
LPTS + VAR 9.10 0.68 0.81 0.73 0.56 3.97%
R-SPARSE (64) 9.03 0.79 0.76 0.75 0.64 -2.55%
R-SPARSE (128) 9.12 0.77 0.77 0.75 0.63 -1.51%
   Gemma3-4B-Instruct 17.29 0.72 0.84 0.72 0.62 -
ACT 35.62 0.65 0.76 0.70 0.51 9.94%
WT 421.95 0.35 0.44 0.58 0.49 34.86%
D-PTS 35.94 0.70 0.76 0.70 0.60 4.58%
S-PTS 35.84 0.71 0.77 0.70 0.60 3.93%
VAR 33.25 0.60 0.76 0.63 0.54 5.04%
CLACT 39.22 0.66 0.74 0.67 0.59 8.01%
Amber-Pruner 35.56 0.67 0.76 0.68 0.61 5.91%
LPTS 19.55 0.65 0.73 0.70 0.55 9.19%
LPTS + VAR 19.13 0.65 0.74 0.71 0.53 9.82%
R-SPARSE (64) 17.04 0.69 0.76 0.69 0.60 5.17%
R-SPARSE (128) 16.17 0.68 0.75 0.70 0.61 5.16%
Table 12: Semi-Structured 8:16 Sparsification - performance Metrics, for calibration, when it is required, and perplexity we use WikiText2. Average Drop is computed without accounting for perplexity.
Pruning WikiText2 $\downarrow$ ARC Easy BoolQ PIQA WinoGrande Average Drop %
   Llama2-7B-chat 6.94 0.74 0.80 0.76 0.66 -
ACT 8.12 0.69 0.75 0.73 0.63 5.37%
WT 20.47 0.64 0.76 0.72 0.61 7.84%
D-PTS 6.92 0.70 0.73 0.75 0.64 4.63%
S-PTS 6.93 0.70 0.73 0.75 0.66 3.87%
VAR 6.67 0.69 0.72 0.75 0.65 4.85%
CLACT 6.54 0.71 0.74 0.75 0.64 3.98%
CLACT + PTS 7.00 0.69 0.72 0.73 0.64 5.63%
CLACT + VAR 6.72 0.69 0.73 0.75 0.64 5.07%
R-SPARSE (64) 7.75 0.69 0.71 0.73 0.64 5.91%
R-SPARSE (128) 7.82 0.68 0.69 0.74 0.61 7.93%
Amber-Pruner 8.10 0.66 0.75 0.73 0.66 5.32%
Amber-Pruner + PTS 6.90 0.68 0.72 0.72 0.65 6.16%
Amber-Pruner + VAR 6.66 0.70 0.72 0.74 0.65 4.74%
LPTS 7.50 0.69 0.66 0.74 0.63 8.15%
LPTS + VAR 7.52 0.69 0.67 0.74 0.64 6.87%
   Llama3.1-8B-Instruct 7.21 0.82 0.84 0.80 0.73 -
ACT 10.32 0.75 0.80 0.76 0.66 7.38%
WT 22.56 0.51 0.64 0.63 0.54 27.26%
D-PTS 10.34 0.76 0.80 0.76 0.66 6.79%
S-PTS 10.31 0.76 0.80 0.75 0.66 7.30%
VAR 10.67 0.74 0.79 0.75 0.66 8.30%
CLACT 10.67 0.73 0.79 0.74 0.66 8.60%
CLACT + PTS 10.68 0.74 0.79 0.74 0.65 8.55%
CLACT + VAR 10.15 0.74 0.79 0.75 0.64 8.59%
R-SPARSE (64) 11.42 0.75 0.77 0.75 0.66 8.44%
R-SPARSE (128) 10.43 0.75 0.78 0.74 0.66 8.49%
Amber-Pruner 10.16 0.73 0.80 0.75 0.68 7.13%
Amber-Pruner + PTS 10.17 0.75 0.80 0.75 0.66 7.42%
Amber-Pruner + VAR 9.94 0.74 0.80 0.75 0.64 8.39%
LPTS 10.04 0.76 0.79 0.77 0.65 7.19%
LPTS + VAR 10.26 0.77 0.78 0.76 0.66 7.15%
   Qwen2.5-7B-Instruct 7.46 0.69 0.86 0.75 0.60 -
ACT 8.61 0.66 0.87 0.73 0.53 4.38%
WT 40.79 0.59 0.82 0.67 0.52 9.54%
D-PTS 8.61 0.80 0.87 0.77 0.68 -8.28%
S-PTS 8.84 0.80 0.86 0.76 0.67 -7.24%
VAR 11.91 0.69 0.72 0.75 0.65 1.93%
CLACT 8.94 0.77 0.85 0.73 0.65 -4.02%
CLACT + PTS 8.94 0.77 0.86 0.83 0.67 -5.06%
CLACT + VAR 8.87 0.76 0.84 0.71 0.65 -2.90%
R-SPARSE (64) 8.12 0.82 0.79 0.77 0.69 -6.90%
R-SPARSE (128) 8.24 0.80 0.79 0.77 0.67 -5.40%
Amber-Pruner 8.80 0.77 0.86 0.74 0.69 -6.20%
Amber-Pruner + PTS 8.79 0.77 0.85 0.73 0.64 -3.40%
Amber-Pruner + VAR 8.73 0.75 0.85 0.74 0.65 -3.60%
LPTS 8.23 0.69 0.83 0.75 0.57 1.70%
LPTS + VAR 8.21 0.70 0.83 0.73 0.56 2.80%
   Gemma3-4B-Instruct 17.29 0.72 0.84 0.72 0.62 -
ACT 25.31 0.70 0.81 0.71 0.55 4.76%
WT 198.53 0.39 0.60 0.62 0.52 26.11%
D-PTS 25.17 0.70 0.81 0.71 0.54 5.16%
S-PTS 25.40 0.75 0.82 0.74 0.63 -1.54%
VAR 23.93 0.75 0.81 0.73 0.65 -1.87%
CLACT 25.85 0.75 0.81 0.71 0.61 0.60%
CLACT + PTS 26.03 0.73 0.81 0.70 0.63 0.50%
CLACT + VAR 24.78 0.74 0.81 0.70 0.63 0.54%
R-SPARSE (64) 15.39 0.76 0.80 0.74 0.64 -1.36%
R-SPARSE (128) 14.55 0.74 0.80 0.73 0.63 -0.44%
Amber-Pruner 25.11 0.74 0.82 0.70 0.63 0.08%
Amber-Pruner + PTS 25.28 0.75 0.81 0.70 0.63 0.17%
Amber-Pruner + VAR 23.97 0.74 0.81 0.71 0.64 -0.16%
LPTS 15.73 0.70 0.79 0.73 0.56 4.21%
LPTS + VAR 15.68 0.71 0.79 0.72 0.57 3.41%
Table 13: Llama3.1-8B-Instruct with 8:16 activation sparsity. LS+L-PTS indicates Learnable Diagonal Scale + Learnable Shift, “Layers” indicates the subset of linear layers where the method was applied. Drop is computed without accounting for perplexity.
Method Layers PPL BoolQ WinoGrande PIQA ARC Easy ARC Chal. HellaSwag OpenBookQA RTE MMLU Lambada stand. Lambada (OpenAI) Drop % $\downarrow$
ORIGINAL 0.8391 0.7340 0.8003 0.8207 0.5196 0.5905 0.3420 0.6859 0.6790 0.6569 0.7308
LS+L-PTS all 9.6036 0.7841 0.6638 0.7715 0.7647 0.4514 0.5294 0.2720 0.6318 0.5521 0.5686 0.6625 10.90%
LS+L-PTS k,o,gate,down 8.3483 0.8205 0.7111 0.7889 0.7887 0.4659 0.5591 0.3200 0.6462 0.6060 0.6123 0.7046 5.43%
LS+L-PTS k,v,gate,down 8.0821 0.8352 0.7174 0.7867 0.7992 0.4898 0.5651 0.3260 0.6643 0.6322 0.6262 0.7108 3.56%
LS+L-PTS + VAR all 9.4983 0.7872 0.6606 0.7606 0.7601 0.4334 0.5372 0.2880 0.6390 0.5532 0.5729 0.6689 10.60%
LS+L-PTS + VAR k,o,gate,down 8.2930 0.8116 0.7135 0.7851 0.7908 0.4838 0.5634 0.3300 0.6498 0.6095 0.6189 0.7079 4.64%
LS+L-PTS + VAR k,v,gate,down 8.0259 0.8306 0.7269 0.7851 0.7955 0.4863 0.5673 0.3260 0.6715 0.6327 0.6317 0.7143 3.36%